A machine cost $15,000. Consider the machine capital recover cost is about $5,500 at an interest rate of 6%. What will be the expected salvage value after 6 years down the road?
Answer #1

Let salvage value be S, then

Capital recovery amount = 15000 * (A/P, 6%,6) - S *(A/F, 6%,6)

As per condition given in question

5500 = 15000 * (A/P, 6%,6) - S *(A/F, 6%,6)

5500 = 15000 * 0.203363 - S * 0.143363

S * 0.143363 = 15000 * 0.203363 - 5500 = -2449.56

S = -2449.56 / 0.143363 = -17086.46

It is additional cost to dispose the equipment at the end of 6 yrs